# SenseFlow: Scaling Distribution Matching for Flow-based Text-to-Image Distillation

### v1

@@ A.5 More Experimental Results
Results of Different Backbone Scales.
We evaluate the impact of different VFM backbone scales (ViT-S, B, and L) in the discriminator on SDXL distillation. Interestingly, the results (Tab.1) do not follow a monotonic trend with respect to model size. ViT-B achieves the best FID-T, while ViT-S yields higher CLIP Score and ImageReward. ViT-L slightly outperforms others on HPSv2 and PickScore. These findings suggest that different backbone scales offer different trade-offs in semantic alignment versus visual fidelity, and that larger backbones do not necessarily guarantee consistent improvements across all metrics. This observation is partially consistent with findings in the ADDsauer2024adversarial paper, which also noted diminishing returns when scaling the discriminator. In our main paper, we adopt ViT-L as the default backbone for the VFM-based discriminator.
@@
Examples from T2I-CompBench.
As shown in Fig. 8, we present visual comparisons of different methods on SDXL using the T2I-CompBench benchmark. These qualitative results clearly highlight the superiority of our approach across multiple aspects, including color fidelity (rows 1 and 2), shape consistency (row 3), material and texture (row 4), and complex spatial arrangements (row 5). Additionally, we also present more examples of our method on SDXL in Fig. 9.
Figure 8: Examples from T2I-CompBench.
@@
  • A stylized digital art poster with the word "SenseFlow" written in flowing smoke from a stage spotlight
@@
  • cartoon dog sits at a table, coffee mug on hand, as a room goes up in flames. "Help" the dog is yelling
@@
A.7 Licenses for existing assets
We use only publicly available and properly licensed open-source datasets and pretrained models in this work. All assets are cited in the main paper, and their licenses explicitly permit academic usage, redistribution, or derivative works under specific conditions. Below is a list of the key assets used and their associated licenses:
  • LAION-5B: Licensed under CC-BY 4.0.
    A large-scale text-image dataset used in pretraining and evaluation contexts.
  • COCO-2017: Licensed under a custom non-commercial research license.
    Commonly used for generation evaluation.
  • Stable Diffusion XL: Licensed under CreativeML Open RAIL++-M.
    Used as a diffusion based teacher model in our distillation framework.
  • Stable Diffusion 3.5: Licensed under CreativeML Open RAIL++-M.
    Used as a large flow-matching base model.
  • FLUX.1-dev: Licensed under CreativeML Open RAIL++-M.
    Used as a large flow-matching base model.
  • DINOv2: Licensed under Apache 2.0.
    Used as the frozen vision foundation backbone in our discriminator design.
  • OpenCLIP: Licensed under Apache 2.0.
    Serves as the text encoder for prompt conditioning in the discriminator.
  • T2I-CompBench: Licensed under the MIT License.
    Used for benchmark comparison of compositional generation performance.
All assets were used in accordance with their respective licenses.

### v2

@@ B.8 Comparison to DMD2 and detailed component ablations
B.9 Qualitative comparisons
We further provide qualitative side-by-side comparisons for SDXL, SD 3.5 Large, and FLUX. For each model family, we show samples from SenseFlow and strong baselines under the same prompts. As illustrated in Figures 1315, SenseFlow tends to produce images with sharper details, cleaner structure, and more faithful overall quality.
Figure 13: Qualitative SDXL comparisons under shared prompts. Each row corresponds to one prompt, showing SenseFlow and strong SDXL baselines side by side.
Figure 14: Qualitative SD 3.5 Large comparisons under shared prompts. SenseFlow produces visually appealing images with improved composition and text–image consistency compared to SD 3.5-Large-Turbo and SD 3.5-Large.
Figure 15: Qualitative FLUX comparisons under shared prompts. SenseFlow provides clearer structures and better alignment with the textual descriptions than Hyper-FLUX and FLUX-Turbo-Alpha.
B.10 More Ablation Study Results and Visualization Samples
Effect of Different Adversarial Loss Weights. In our main experiments, the hyperparameter $\lambda_{\text{G}}$ in Algorithm 1, Line 22, is set to 0.5, 0.1, and 2.0 for SDXL, SD 3.5 Large, and FLUX.1 dev, respectively. To further investigate the impact of this hyperparameter, we conduct an ablation study using SDXL as an example, decreasing $\lambda_{\text{G}}$ to 0.25. The results are presented in Tab. 10. We observe that setting $\lambda_{\text{G}}=0.5$ leads to improved performance across most metrics, including CLIP Score, HPSv2, PickScore, and ImageReward. Notably, this configuration achieves the best scores on HPSv2, PickScore, and ImageReward among all methods in Tab. 1. These results highlight the strong semantic and visual supervision capabilities of our VFM-based discriminator.
Results of Different Backbone Scales. We evaluate the impact of different VFM backbone scales (ViT-S, B, and L) in the discriminator on SDXL distillation. Interestingly, the results (Tab.1) do not follow a monotonic trend with respect to model size. ViT-B achieves the best FID-T, while ViT-S yields higher CLIP Score and ImageReward. ViT-L slightly outperforms others on HPSv2 and PickScore. These findings suggest that different backbone scales offer different trade-offs in semantic alignment versus visual fidelity, and that larger backbones do not necessarily guarantee consistent improvements across all metrics. This observation is partially consistent with findings in the ADD(Sauer et al., 2024b) paper, which also noted diminishing returns when scaling the discriminator. In our main paper, we adopt ViT-L as the default backbone for the VFM-based discriminator.
More visualization results of our methods on SD 3.5 Large and SDXL. As shown in Fig. 16 and Fig. 17, we present more samples produced by out 4-step generator distilled from SD 3.5 Large and SDXL, separately. The prompts of these samples are listed later in our appendix.
@@ B.10 More Ablation Study Results and Visualization Samples
  • A stylized digital art poster with the word ”SenseFlow” written in flowing smoke from a stage spotlight
@@
  • cartoon dog sits at a table, coffee mug on hand, as a room goes up in flames. ”Help” the dog is yelling